A sample of 66 observations will be taken from an infinite population. The population proportion equals 0.12. The probability that the sample proportion will be less than 0.1969 is
Answer #1

From the given information

n=66

P=population proportion =0.12

p is sample proportion

We know that sample proportion is approximately normally distributed with

Mean=P = 0.12

Standard deviation = S.d.=​​​​​​

=​​​​​​

=0.04

Probability that sample proportion is less than 0.1969

=p( p < 0.1969)

z=(p-mean) /s.d. ~ N(0,1)

=p( z < (0.1969-0.12)/0.04)

=p(z < 1.92)

= 0.9726